The length of a rectangle is 4 cm longer than the width. If the perimeter is 20 cm, what is the area of the rectangle?
Answer by checkley77(12569)

L=W+4
P=2L+2W
20=2(W+4)2W
20=2W+8+2W
20=4W+8
4W=20-8
4W=12
W=12/4
W=3 ANS. FOR THE WIDTH.
L=3+4=7 ANS. FOR THE LENGTH.
PROOF:
20=2*7=2*3
20=14+6
20=20